OpenStudy (anonymous):

COMPLEX ANALYSIS @dan815 @jagatuba

OpenStudy (anonymous):

Let f be analytic inside and on the simple closed contour Gamma. What is the value of \[\frac{ 1 }{ 2\pi*i }\int\limits_{\Gamma}^{}\frac{ f(z) }{ z-z_o }dz\] when z_o lies outside Gamma?

terenzreignz (terenzreignz):

Shouldn't it simply be zero?

OpenStudy (anonymous):

Yep :)

terenzreignz (terenzreignz):

Then what was the point of this? -_-

OpenStudy (anonymous):

lolz